Shift report 20011005 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Oct 05 2001 - 15:58:48 EDT

  • Next message: Flemming Videbaek: "RHIC - Experimental Maintenance Shutdown next week"

    ------------------------------------------------------------
    Date: Fri Oct  5 15:58:47 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#9 pages 96 - 97
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam hours :  8:32-11:13, 12:27-16:00
                      about 6 hours...
    
    
    2.  MRS moved 35-degree, set 4A-polarity (700amp)
        FFS 12-degree, BFS 3-degree, 1/5 B-polarity 
    
        Data 5165, 5166 : MRS (4A), FS (Magnet OFF)
                          total ~116k trigger 1       
        
        Data 5167, 5174, 5176, 5178 : MRS (4A), FS (1/5B)
                                      total ~ 0.6M trigger 1
    
    3.  No new problems detected.
    
        Alarms "T3-map" on DAQ window : several times RED, 
                       after few seconds, RED disappeared.
        Print job from pii5 : not available.
         -> need to reboot (Betty's suggestion)
         -> root files from brntuple are in ~operator/ntuple/data
         -> no print-output from brntuple since midnight-shift
        TPM2 figures from "XTPI" : looks problems.
         -> Ramiro knows this fact.      
                            
                               
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Oct 05 2001 - 15:59:24 EDT